OpenStudy (brinazarski):

Solve for x: Arctan(x/2) + Arctan(2x/3) = Pi/4

OpenStudy (tkhunny):

Do you remember this one? \(arctan(A) + arctan(B) = arctan\left(\dfrac{A + B}{1 - AB}\right) \)

OpenStudy (brinazarski):

Well that helps. Thank you!

OpenStudy (brinazarski):

I didn't know that was the formula lol

OpenStudy (brinazarski):

... okay, I'm still getting it wrong.

OpenStudy (anonymous):

i am beginning to hate the arctan @wio

OpenStudy (tkhunny):

What did you get for \(\dfrac{A+B}{1-AB}\)?